if a = 15 80° and b = 12i + 16j what is the magnitude of a - b


1
Expert's answer
2021-05-17T12:10:02-0400

"\vec{A}=(15,80^{\circ})"

     "=15cos80^{\circ}+15sin80^{\circ}"

     "=2.61i+14.77j"


"\vec{B}=12i-16j"


SO, Vector "({A-B}=(2.61i+14.77j)-(12i-16j)"

       "=-9.39i+30.77j"


Magnitude "|(A-B|=\sqrt{(-9.39)^2+(30.77)^2}=\sqrt{88.17+946.79}=\sqrt{1034.96}=32.17"